# Aligning Machiavellian Agents: Behavior Steering via Test-Time Policy Shaping

### v1

@@ Abstract
The deployment of decision-making AI agents presents a critical challenge in maintaining alignment with human values or guidelines while operating in complex, dynamic environments. Agents trained solely to achieve their objectives may adopt harmful behavior, exposing a key trade-off between maximizing the reward function and maintaining the alignment. For the pre-trained agents, ensuring alignment is particularly challenging, as retraining can be a costly and slow process. This is further complicated by the diverse and potentially conflicting attributes representing the ethical values for alignment. To address these challenges, we propose a test-time alignment technique based on model-guided policy shaping. Our method allows precise control over individual behavioral attributes, generalizes across diverse reinforcement learning (RL) environments, and facilitates a principled trade-off between ethical alignment and reward maximization without requiring agent retraining. We evaluate our approach using the MACHIAVELLI benchmark, which comprises 134 text-based game environments and thousands of annotated scenarios involving ethical decisions. The RL agents are first trained to maximize the reward in their respective games. At test time, we apply policy shaping via scenario-action attribute classifiers to ensure decision alignment with ethical attributes. We compare our approach against prior training-time methods and general-purpose agents, as well as study several types of ethical violations and power-seeking behavior. Our results demonstrate that test-time policy shaping provides an effective and scalable solution for mitigating unethical behavior across diverse environments and alignment attributes.
@@ 4.2 Policy Shaping with Agent Interpolation
Given a set of $N$ attributes $\{(k_{i},v_{i})\}_{i=1}^{N},$ where $k_{i}$ denotes the attribute and $v_{i}\in\{0,1\}$ indicates preference (0 for minimization, 1 for maximization), the classifier action probabilities for the action space $\mathcal{A}$ are computed as:
@@ A.1 Definitions of Attributes
The MACHIAVELLI benchmark defines thirteen ethical violations, four power-seeking attributes, and one disutiltiy attribute, as listed in Table 2 with corresponding definitions. For our study, we focus on a subset of these attributes and select ten test games from the full set of thirty. The criteria and procedure for selecting these subsets are described in the following section.
@@ A.2 Method for Selection of Games and Attributes
To select a subset of ethical violations, we identify the ten most frequently occurring attributes across MACHIAVELLI game scenarios that involve choice-based decision making. Since each game consists of multiple scenarios, many of which do not require the agent to make decisions, we focus specifically on choice-based scenarios where ethical optimization is relevant. The distribution of attributes in these scenarios is shown in Fig. 8. Based on this distribution, we select the top ten violations, excluding “other”, “promises”, and “vandalism”. We also include all power and utility attributes, as they appear consistently across all games.
@@ C.1 Classifier Training
power.money 0.890 0.891 0.453
power.physical 0.893 0.850 0.192
power.social 0.903 0.866 0.253
power.utility 0.866 0.892 0.271
disutility 0.878 0.914 0.221
@@ C.2 Classifier Results
As noted in Section 4.1, the number of positive choice scenarios for each attribute is severely imbalanced compared to the number of negative cases. Attributes such as “killing” and “trespassing” have nearly 100 positive examples, while the number of negative examples approaches 20,000. During training, we counteract this imbalance by randomly sampling, without replacement, an equal number of positive and negative examples for each attribute classifier model. However, this imbalance affects both training and evaluation, since fewer examples are available overall for attributes such as “fairness”, resulting in lower precision scores when classifying these attributes. This is observed in Table 3.
@@ Appendix D Extended Policy Shaping Results
Figure 10: Alignment results on the top five ethical violations and total power, of the RL agent, RL agent with an artificial conscience (AC), LLM agent, and the Good LLM agent. A more harmful agent will have a larger area.
Baseline results for RL, RL-AC, and LLM agents are presented in Section 5 of the main text and illustrated in Figure 10. Building on these findings, Table 4 reports outcomes across all $\alpha$ values for RL and RL-AC agents, showing the effect of policy shaping with our attribute classifiers aimed at reducing unethical behavior. We also attempt to further improve the RL-AC agent’s performance. These trends are illustrated in the Pareto front plots for the top five attributes in the main text, and the other five attributes in Figure 11. As $\alpha$ increases, the total number of ethical violations decreases, with $\alpha=0.8$ yielding the lowest violation count. This approach outperforms policy shaping applied to the RL-AC agent, which showed only modest improvements in attributes such as stealing, trespassing, and physical harm. The limited effect is likely due to the RL-AC agent’s training process already reducing unethical behavior and possibly causing a distribution shift in its policy.
@@ D.1 Reversing Training-Time Alignment
Similarly, we present outcomes for additional $\alpha$ values on RL and RL-AC agents when intentionally steering them in the opposite direction of their training-time alignment, effectively reversing their learned alignment behavior. We examine this in the context of ethical violations, power-seeking actions, and disutility. These results are presented in Table 5, and also illustrated in the Pareto front in Figure 7 of the main text for the top five attributes, and Figure 12 for the remaining five attributes. In this analysis, we also examine the effects of steering the base RL agent further in the opposite direction to study the impact of steering in both directions. We observe similar results to those presented in Sec. 5.4, where the RL-AC exhibits an increase in unethical behavior. However, for the RL agent, the increase in violations is less pronounced, likely due to the already high number of unethical actions taken during gameplay, as well as potential false positives from attribute predictors skewing the results. This pattern is also observed with some ethical violations in the RL-AC, such as fairness, which has one of the lowest accuracy scores.
@@ D.2 Classifier Accuracy Impact on Policy Shaping
Observing the results in Table 4 and Figure 11, a few attributes, such as “fairness” and “killing,” exhibit a weaker Pareto front, likely due to inaccuracies in their attribute classifications stemming from the smaller number of positive training examples available for these attributes. A similar pattern appears in Figure 12, where reversing training-time alignment for the same attributes produces a partially weakened effect because of classifier limitations, though it still leads to an overall improvement in alignment. Given the nuanced nature of defining ethical attributes through human annotation, such effects may vary across domains and under distributional shifts between training and test data. With MACHIAVELLI, classifiers were explicitly evaluated under distribution shift, using distinct text games that differed substantially in content between training and testing. However, in other application domains, ensuring sufficient variation between training and test data will be important to mitigate the impact of attribute classifier distribution shift.
@@ D.3 Statistical Significance of Results
We present the results of statistical significance for the RL-$\alpha 0.5$ and RL-$\alpha 1.0$ agents, compared to the baseline RL agent, across the ten test games. Table 6 reports the mean and standard deviation of scores for each attribute. Statistical significance was assessed using the Wilcoxon signed-rank test for non-normally distributed variables, with normality evaluated using the Shapiro-Wilk test. The results show that, for most attributes, the improvements of the RL-$\alpha$ agents over the baseline are statistically significant. However, attributes “Money”, “Stealing”, “Spying” and “Trespassing,” were not statistically significant likely due to the high variability of these attributes in the baseline RL agent and the limited number of independent games. Nevertheless, the mean scores for all attributes are consistently lower for the RL-$\alpha$ agents, indicating systematic improvements. We also note that the standard deviation across attributes for the RL-$\alpha$ agents is consistently lower than that of the base RL agent. This is likely due to the nature of the games, where points are tied to unethical behaviors, leading to high variability in the baseline agent’s scores across traits, which can differ substantially between games.
@@ Appendix E Attribute Correlations
Following the results presented in Sec. 5.3 and Fig. 6 of the main text, we further analyze correlations between attributes used in policy shaping. When agents are shaped to maximize certain attributes, stronger negative correlations are observed between some attributes, such as “spying” and “deception” relative to “killing”, than when shaping policies to minimize attributes. This likely reflects inherent trade-offs between these behaviors within game contexts. Additionally, attributes with fewer occurrences across games, such as “fairness” or “stealing,” do not exhibit weaker correlations than more frequently occurring attributes like “deception” or “manipulation,” suggesting that attribute frequency alone does not determine correlation strength. These results indicate that correlations among ethical attributes should be considered when selecting which attributes to emphasize during alignment, as targeting highly correlated attributes may amplify or offset specific behaviors.
@@ Appendix F Agent Trajectory Viewer
To facilitate debugging and analyze trends in agent behavior across games, we developed a Python-based trajectory viewer module that visualizes agent paths through scenarios, their choices at each stage, and the ethical attributes associated with those decisions. This module is included in our code repository, and examples of generated trajectories are shown in Figure 13, with a close-up provided in Figure 14. Nodes are highlighted to indicate the occurrence of ethical violations, and hovering over a node reveals the scenario text, available actions, and associated attributes (shown in Fig. 14. This tool is particularly useful for identifying situations in which agents are compelled to select an unethical action (e.g., when all available options involve an ethical violation), or for detecting loops between nodes that result in longer or shorter trajectories, thereby affecting the total number of points earned.
@@ Appendix G Multi-Attribute Alignment
Another focus of our alignment study is pluralistic alignment, which refers to optimizing for more than one attribute simultaneously. We evaluate the effectiveness of our policy-shaping approach in achieving alignment across multiple attributes and examine how varying the target level of these attributes (e.g., increasing or decreasing them) influences RL agents. For this analysis, we consider the three most frequently occurring attributes in MACHIAVELLI that do not have strong correlations: physical harm, deception, and non-physical harm. Baseline results for the Oracle Agent across all combinations of these attributes are presented in Table 7. Similarly, results for policy shaping applied to the RL agent are reported in Table 7. We also target variations of reduction of one ethical violation, and increasing another, assuming equal weighing between attributes.
Looking at Table 1 in the main text, and comparing our policy-shaped agents to the baseline RL results, we observe a consistent trend with prior experiments: targeting the reduction of an attribute leads to fewer violations, while targeting an increase results in higher violation counts. However, we also observe high standard deviations for the first target attribute, likely due to interactions such as correlations or inverse correlations between attributes when multiple targets are aligned. When minimizing both attributes, we observe an overall reduction in total violations. These results highlight the importance of understanding relationships between attributes and selecting appropriate weights when aligning to multiple ethical objectives.
Table 7: Comparison of RLRL-$\alpha 0.5$ and Oracle agent performance across three ethical violation types, under dual-target conditions $(v_{0},v_{1})$. Each column corresponds to the first targeted attribute $v_{0}$, while row groups reflect the second targeted attribute $v_{1}$. Metrics include targeted attributes, total violations, and game reward (i.e., points and achievements). Metrics not directly optimized by an agent (e.g., game points or cases where attributes $v_{0}$ and $v_{1}$ are equal and not targeted) are aggregated, with mean and standard deviation across agents reported.
@@ Appendix H LLM Agent Prompts
Following the MACHIAVELLI framework, we reproduce their four agents using LLAMA 2 7B. Two of these agents, the “LLM Agent” and the “Good LLM Agent”, are discussed in Section 3 of the main text. In addition, we introduce two new agents that incorporate chain-of-thought (CoT) prompting: the “CoT Agent” and the “Good CoT Agent.” The latter includes an ethical prompt similar to that used for the “Good LLM Agent.” CoT agents prompt the LLM twice per scenario: first to generate reasoning for each action, and then to select a final action. Results for all agents are reported in Section 8.

### v2

@@ Abstract
The deployment of decision-making AI agents presents a critical challenge in maintaining alignment with human values or guidelines while operating in complex, dynamic environments. Agents trained solely to achieve their objectives may adopt harmful behavior, exposing a key trade-off between maximizing the reward function and maintaining alignment. For pre-trained agents, ensuring alignment is particularly challenging, as retraining can be a costly and slow process. This is further complicated by the diverse and potentially conflicting attributes representing the ethical values for alignment. To address these challenges, we propose a test-time alignment technique based on model-guided policy shaping. Our method allows precise control over individual behavioral attributes, generalizes across diverse reinforcement learning (RL) environments, and facilitates a principled trade-off between ethical alignment and reward maximization without requiring agent retraining. We evaluate our approach using the MACHIAVELLI benchmark, which comprises 134 text-based game environments and thousands of annotated scenarios involving ethical decisions. The RL agents are first trained to maximize the reward in their respective games. At test time, we apply policy shaping via scenario-action attribute classifiers to ensure decision alignment with ethical attributes. We compare our approach against prior training-time methods and general-purpose agents, as well as study several types of ethical violations and power-seeking behavior. Our results demonstrate that test-time policy shaping provides an effective and scalable solution for mitigating unethical behavior across diverse environments and alignment attributes.
@@ 4.2 Policy Shaping with Agent Interpolation
Given a set of NN attributes $\{(k_{i},v_{i})\}_{i=1}^{N},$ where $k_{i}$ denotes the attribute and $v_{i}\in\{0,1\}$ indicates preference (0 for minimization, 1 for maximization), the classifier action probabilities for the action space $\mathcal{A}$ are computed as:
@@ A.1 Definitions of Attributes
The MACHIAVELLI benchmark defines thirteen ethical violations, four power-seeking attributes, and one disutility attribute, as listed in Table 2, with corresponding definitions. For our study, we focus on a subset of these attributes and select ten test games from the full set of thirty. The criteria and procedure for selecting these subsets are described in the following section.
@@ A.2 Method for Selection of Games and Attributes
To select a subset of ethical violations, we identify the ten most frequently occurring attributes across MACHIAVELLI game scenarios that involve choice-based decision making. Since each game consists of multiple scenarios, many of which do not require the agent to make decisions, we focus specifically on choice-based scenarios where ethical optimization is relevant. The distribution of attributes in these scenarios is shown in Fig. 8. Based on this distribution, we select the top ten violations, excluding “other,” “promises,” and “vandalism”. We also include all power attributes and disutility, as they appear consistently across all games.
@@ C.2 Classifier Results
As noted in Section 4.1, the number of positive choice scenarios for each attribute is severely imbalanced compared to the number of negative cases. Attributes such as “killing” and “trespassing” have nearly 100 positive examples, while the number of negative examples approaches 20,000. During training, we counteract this imbalance by randomly sampling, without replacement, an equal number of positive and negative examples for each attribute classifier model. However, this imbalance affects both training and evaluation, since fewer examples are available overall for attributes such as “fairness,” resulting in lower precision scores when classifying these attributes. This is observed in Table 3.
@@ Appendix D Extended Policy Shaping Results
Figure 10: Alignment results on the top five ethical violations and total power of the RL agent, RL agent with an artificial conscience (AC), LLM agent, and the Good LLM agent. A more harmful agent will have a larger area.
Baseline results for RL, RL-AC, and LLM agents are presented in Section 5 of the main text and illustrated in Figure 10. Building on these findings, Table 4 reports outcomes across all $\alpha$ values for RL and RL-AC agents, showing the effect of policy shaping with our attribute classifiers aimed at reducing unethical behavior. We also attempt to further improve the RL-AC agent’s performance. These trends are illustrated in the Pareto front plots for the top five attributes in the main text, and the other five attributes in Figure 11. As $\alpha$ increases, the total number of ethical violations decreases, with $\alpha=0.8$ yielding the lowest violation count. This approach outperforms policy shaping applied to the RL-AC agent, which showed only modest improvements in attributes such as “stealing,” “trespassing,” and “physical harm”. The limited effect is likely due to the RL-AC agent’s training process already reducing unethical behavior and possibly causing a distribution shift in its policy.
@@ D.1 Reversing Training-Time Alignment
Similarly, we present outcomes for additional $\alpha$ values on RL and RL-AC agents when intentionally steering them in the opposite direction of their training-time alignment, effectively reversing their learned alignment behavior. We examine this in the context of ethical violations, power-seeking actions, and disutility. These results are presented in Table 5, and also illustrated in the Pareto front in Figure 7 of the main text for the top five attributes, and Figure 12 for the remaining five attributes. In this analysis, we also examine the effects of steering the base RL agent further in the opposite direction to study the impact of steering in both directions. We observe similar results to those presented in Sec. 5.4, where the RL-AC exhibits an increase in unethical behavior. However, for the RL agent, the increase in violations is less pronounced, likely due to the already high number of unethical actions taken during gameplay, as well as potential false positives from attribute predictors skewing the results. This pattern is also observed with some ethical violations in the RL-AC agent, such as fairness, which has one of the lowest accuracy scores.
@@ D.2 Classifier Accuracy Impact on Policy Shaping
Observing the results in Table 4 and Figure 11, some attributes, such as “fairness” and “killing,” exhibit weaker Pareto fronts. This is likely due to inaccuracies in their attribute classifiers, which stem from the relatively small number of positive training examples for these attributes. A similar pattern appears in Figure 12: reversing training-time alignment for the same attributes yields only a partially weakened effect because of these classifier limitations, although it still produces an overall improvement in alignment.
Given the nuanced nature of defining ethical attributes through human annotation, such effects may vary across domains and under distributional shifts between training and test data. In MACHIAVELLI, classifiers were explicitly evaluated under distribution shift by using distinct text-game environments that differed substantially in content between training and testing. In other application domains, however, ensuring adequate variation between training and test data will be essential to mitigate the impact of attribute-classifier distribution shift.
@@ D.3 Statistical Significance of Results
We present the results of statistical significance for the RL-$\alpha 0.5$ and RL-$\alpha 1.0$ agents, compared to the baseline RL agent, across the ten test games. Table 6 reports the mean and standard deviation of scores for each attribute. Statistical significance was assessed using the Wilcoxon signed-rank test for non-normally distributed variables, with normality evaluated using the Shapiro-Wilk test. The results show that, for most attributes, the improvements of the RL-$\alpha$ agents over the baseline are statistically significant. However, attributes “money,” “stealing,” “spying,” and “trespassing,” were not statistically significant likely due to the high variability of these attributes in the baseline RL agent and the limited number of independent games. Nevertheless, the mean scores for all attributes are consistently lower for the RL-$\alpha$ agents, indicating systematic improvements. We also note that the standard deviation across attributes for the RL-$\alpha$ agents is consistently lower than that of the base RL agent. This is likely due to the nature of the games, where points are tied to unethical behaviors, leading to high variability in the baseline agent’s scores across traits, which can differ substantially between games.
@@ Appendix E Attribute Correlations
Following the results presented in Sec. 5.3 and Fig. 6 of the main text, we further analyze correlations between attributes used in policy shaping. When agents are shaped to maximize certain attributes, stronger negative correlations are observed between some attributes, such as “spying” and “deception” relative to “killing,” than when shaping policies to minimize attributes. This likely reflects inherent trade-offs between these behaviors within game contexts. Additionally, attributes with fewer occurrences across games, such as “fairness” or “stealing,” do not exhibit weaker correlations than more frequently occurring attributes like “deception” or “manipulation,” suggesting that attribute frequency alone does not determine correlation strength. These results indicate that correlations among ethical attributes should be considered when selecting which attributes to emphasize during alignment, as targeting highly correlated attributes may amplify or offset specific behaviors.
@@ Appendix F Agent Trajectory Viewer
To facilitate debugging and analyze trends in agent behavior across games, we developed a Python-based trajectory viewer module that visualizes agent paths through scenarios, their choices at each stage, and the ethical attributes associated with those decisions. This module is included in our code repository, and examples of generated trajectories are shown in Figure 13, with a close-up provided in Figure 14. Nodes are highlighted to indicate the occurrence of ethical violations, and hovering over a node reveals the scenario text, available actions, and associated attributes (shown in Fig. 14). This tool is particularly useful for identifying situations in which agents are compelled to select an unethical action (e.g., when all available options involve an ethical violation), or for detecting loops between nodes that result in longer or shorter trajectories, thereby affecting the total number of points earned.
@@ Appendix G Multi-Attribute Alignment
Another aspect of our alignment study is pluralistic alignment, which refers to optimizing for more than one attribute simultaneously. We evaluate the effectiveness of our policy-shaping approach in achieving alignment across multiple attributes and examine how varying the target level of these attributes (e.g., increasing or decreasing them) influences RL agents. For this analysis, we consider the three most frequently occurring attributes in MACHIAVELLI that do not have strong correlations: “physical harm,” “deception,” and “non-physical harm”. Baseline results for the Oracle agent across all combinations of these attributes are presented in Table 7, and the corresponding results for policy shaping applied to the RL agent are also shown in Table 7. We additionally examine settings in which one ethical violation is reduced while another is increased, assuming equal weighting across attributes.
Examining Table 1 in the main text and comparing our policy-shaped agents to the baseline RL results, we observe a consistent trend with prior experiments: targeting the reduction of an attribute leads to fewer violations, while targeting an increase results in higher violation counts. However, we also find high standard deviations for the first target attribute, likely due to interactions such as correlations or inverse correlations—between attributes when multiple targets are optimized simultaneously. When minimizing both attributes, we observe an overall reduction in total violations. These findings underscore the importance of understanding inter-attribute relationships and selecting appropriate weights when aligning agents to multiple ethical objectives.
Table 7: Comparison of RL-$\alpha 0.5$ and Oracle agent performance across three ethical violation types, under dual-target conditions $(v_{0},v_{1})$. Each column corresponds to the first targeted attribute $v_{0}$, while row groups reflect the second targeted attribute $v_{1}$. Metrics include targeted attributes, total violations, and game reward (i.e., points and achievements). Metrics not directly optimized by an agent (e.g., game points or cases where attributes $v_{0}$ and $v_{1}$ are equal and not targeted) are aggregated, with mean and standard deviation across agents reported.
@@ Appendix H LLM Agent Prompts
Following the MACHIAVELLI framework, we reproduce their four agents using LLAMA 2 7B. Two of these agents, the “LLM Agent” and the “Good LLM Agent,” are discussed in Section 3 of the main text. In addition, we introduce two new agents that incorporate chain-of-thought (CoT) prompting: the “CoT Agent” and the “Good CoT Agent.” The latter includes an ethical prompt similar to that used for the “Good LLM Agent.” CoT agents prompt the LLM twice per scenario: first to generate reasoning for each action, and then to select a final action. Results for all agents are reported in Section 8.